OpenStudy (anonymous):

Need help...If someone could explain Twenty cards are drawn from a pack of cards at random. The sequence ? is the closest to the probability of drawing a queen or a king from the pack. 5 6 8 2 Q 4 8 3 2 A 2 3 10 J J 4 A 5 3 8 K J 5 8 4 3 6 Q 9 J Q 2 8 Q 9 2 Q A 7 K 3 8 3 2 7 4 J A 6 3 2 A A 6 3 8 6 10 5 4 4 8 6 K J 7 Q 3 A K 2 5 8 9 10 2 4 A 6 7

OpenStudy (anonymous):

@cp9454 can you help me with this one? please

OpenStudy (cp9454):

we have to find probability of getting king and queen from from a deck of 52 cards or the 20 cards that we have driven randomly.

OpenStudy (anonymous):

I need to find which sequence

OpenStudy (anonymous):

There are 52 cards 8 cards are Q & K The sample size is 20 Is this correct so far?

OpenStudy (anonymous):

P(Q or K) = P(Q) = 4/52 1/13 + P(A) = 4/52 1/13 = 1/13 + 1/13 = 2/13 confused on the sequence

OpenStudy (michele_laino):

please, wait a moment

OpenStudy (michele_laino):

here we have to compute the probability associated to each of your four sequences

OpenStudy (michele_laino):

for example what is the probability to get a Q from the first sequence

OpenStudy (michele_laino):

more precisely, what is the probability associated to the first sequence?

OpenStudy (anonymous):

1/52

OpenStudy (anonymous):

4/52

OpenStudy (anonymous):

wait the 5th card out of 20 was a Q

OpenStudy (michele_laino):

I think that the probability for the first sequence, is: \[\frac{1}{{52}} \times \frac{1}{{51}} \times \frac{1}{{50}} \times \frac{1}{{49}} \times \frac{1}{{48}}\]

OpenStudy (michele_laino):

whereas for second sequence, is: \[\frac{1}{{52}}\]

OpenStudy (michele_laino):

for third sequence, is: \[probability = 0\] since there is neither queen nor king

OpenStudy (anonymous):

so the second choice has the best probability

OpenStudy (anonymous):

Did I make this question more difficult than it was?

OpenStudy (michele_laino):

please wait, what is the probability associated to the fourth sequence?

OpenStudy (michele_laino):

we have the king at the fourth place, so the probability is: \[\frac{1}{{52}} \times \frac{1}{{51}} \times \frac{1}{{50}} \times \frac{1}{{49}}\]

OpenStudy (michele_laino):

namely, the second sequence is the right option

OpenStudy (anonymous):

Thank you!